Shallypeace:

The atomic number of an element X is 14, and it's relative atomic mass is 28. How many neutrons does atom of X contain? Write the electronic distribution in an atom of X. Will X react in an electrovalent or convalent manner? Give reasons for your answer

DK4L:

Silicon is a chemical element with the symbol Si, atomic number 14, and relative atomic mass 28. So, in our case X is the Silicon (Si). Silicon-28 is a stable isotope of Silicon with N = A – Z = 28 – 14 = 14 neutrons in the nucleus. The electronic distribution of Silicon is 1s2 2s2 2p6 3s2 3p2.Silicon is a metalloid – the value of his electronegativity is intermediate between metals and non-metals. Also on the outer energy level of the Silicon atom, there are four valent electrons, and it may react with both metals or non-metals to form four covalent bonds and stable electron octet configuration on the outer energy level www.assignmentexpert.com

Shallypeace:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L Silicon is a chemical element with the symbol Si, atomic number 14, and relative atomic mass 28. So, in our case X is the Silicon (Si). Silicon-28 is a stable isotope of Silicon with N = A – Z = 28 – 14 = 14 neutrons in the nucleus. The electronic distribution of Silicon is 1s2 2s2 2p6 3s2 3p2.Silicon is a metalloid – the value of his electronegativity is intermediate between metals and non-metals. Also on the outer energy level of the Silicon atom, there are four valent electrons, and it may react with both metals or non-metals to form four covalent bonds and stable electron octet configuration on the outer energy level www.assignmentexpert.com \(\color{#0cbb34}{\text{End of Quote}}\) Thanks🤗

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shallypeace \(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L Silicon is a chemical element with the symbol Si, atomic number 14, and relative atomic mass 28. So, in our case X is the Silicon (Si). Silicon-28 is a stable isotope of Silicon with N = A – Z = 28 – 14 = 14 neutrons in the nucleus. The electronic distribution of Silicon is 1s2 2s2 2p6 3s2 3p2.Silicon is a metalloid – the value of his electronegativity is intermediate between metals and non-metals. Also on the outer energy level of the Silicon atom, there are four valent electrons, and it may react with both metals or non-metals to form four covalent bonds and stable electron octet configuration on the outer energy level www.assignmentexpert.com \(\color{#0cbb34}{\text{End of Quote}}\) Thanks🤗 \(\color{#0cbb34}{\text{End of Quote}}\) NP